A certain insect has lengths that are normally distributed. The insect has a mean length of 32.4 mm and a standard deviation of 2.5mm. What percent of these insects have a length over 34mm?
Answer #1

Given,

= 32.4, = 2.5

We convert this to standard normal as

P( X < x) = P (Z < x - / )

So,

P( X > 34) = P( Z > 34 - 32.4 / 2.5)

= P( Z > 0.64)

= 0.2611

= 26.11%